Translate the following English sentence into symbol. (3 pts each)

1. No one in this class is wearing pants and a guitarist.

Let:

Domain of x is all persons

A(x): x is wearing pants

B(x): x is a guitarist

C(x): belongs to the class

Answer: __________________________________________________

2. No one in this class is wearing pants and a guitarist.

Let:

Domain of x is persons in this class

A(x): x is wearing pants

B(x): x is a guitarist

Answer: __________________________________________________

3. There is a student at your school who knows C++ but who doesn’t 

know Java.

Let:

Domain: all students at your school

C(x): x knows C++

J(x): x knows Java

Answer: __________________________________________________


1
Expert's answer
2021-05-02T09:43:45-0400

1) No one in this class is wearing pants and a guitarist:

Let:

Domain of x is all persons

A(x): x is wearing pants

B(x): x is a guitarist

C(x): belongs to the class

Answer: "\forall x\left( {C(x) \to \overline {A(x) \wedge B(x)} } \right)"

2) No one in this class is wearing pants and a guitarist.

Let:

Domain of x is persons in this class

A(x): x is wearing pants

B(x): x is a guitarist

Answer: "\forall x\left( {\overline {A(x) \wedge B(x)} } \right)"

3) There is a student at your school who knows C++ but who doesn’t 

know Java.

Let:

Domain: all students at your school

C(x): x knows C++

J(x): x knows Java

Answer: "\exists x\left( {C(x) \wedge \overline {J(x)} } \right)"